News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 17 May 2002 11.00pm EDT

During the last week the observing schedule was replanned to include an observation of the X-ray transient XTE J0929-314 that was accepted as a Target Of Opportunity (TOO) on 10 May. Observations of 0858-771, 1343-601, 0903-573 impacted by the TOO and will be rescheduled in a later load.

Chandra passed through the a 4% lunar eclipse on 11 May with nominal power and thermal performance. The measured temperature decrease of the solar arrays was 2.5 degrees F and compared well with the predicted value of 3 degrees F.

A software patch to increase two glitch counter values was uplinked on May 14 to mitigate a subtle timing condition. The condition has twice caused the EPHIN hardware failure flag to be erroneously set resulting in a loss of science through execution of SCS 107. Both the A- and B-side patches were uplinked.

The observing schedule for the next week is shown below and includes two full orbit observations of Sgr A*.

------------------------------------------
HS 1939+7000           ACIS-S       May 19
Radiation Belts
AQL X-1                ACIS-S
NGC 6482               ACIS-S
1828+487               ACIS-S
HS 1824+6507           ACIS-S
Abell 2124             ACIS-S
RX J1308+2127          ACIS-I
1AXG J110713-1839      ACIS-I       May 21
NGC 5102               ACIS-S
1655-776               ACIS-S
SMC FIELD 7            ACIS-I
SN2001IG               ACIS-S
Radiation Belts                     May 22
Sgr A*                 ACIS-I
Radiation Belts                     May 24
Sgr A*                 ACIS-I       May 25

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports